# MgNO: Efficient Parameterization of Linear Operators via Multigrid

### v1

@@ 5 Experiments
To substantiate the superiority of our method, we conducted comprehensive method comparisons across multiple benchmark scenarios. These benchmarks were categorized based on their association with specific partial differential equations (PDEs, namely Darcy, Navier-Stokes, and Helmholtz benchmarks. Furthermore, we incorporated tasks encompassing both regular and irregular domains, each subject to a variety of boundary conditions. This comprehensive comparison allows us to thoroughly assess the robustness and adaptability of our method, ensuring its effectiveness in real-world scenarios. For clearness, we summarize the benchmarks in Table 5. Please also refer to Section C for more detailed descriptions.
@@
We perform a comprehensive comparison with existing methods: (1) We include the original FNO [23]; (2) UNet [34] and U-NO [32]; (3) multiwavelet neural operator (MWT) [9]; (4) Galerkin transformer (GT) [2]; (5) latent spectral models (LSM) [39]; (6) learned simulator for turbulence (DilResNet) [36]. The baseline models were implemented using their official implementations. For consistency, we executed the baselines with default hyperparameters unless the experiments’ details specified otherwise. In cases where specific experiment details were lacking in the literature, we employed a reasonable network scale to ensure a fair comparison. Regarding our training setup, we adopted a meticulous approach. We trained the baselines using multiple configurations, including their default training settings encompassing different loss functions, training algorithms, and schedulers. We then reported the best results from these trials to mitigate any training-related variations. Notably, our observations revealed that all models benefited from the scheduler utilizing cosine annealing learning rates. Figure 3 will visually demonstrate that all models underwent sufficient training, ensuring that their performance accurately reflects their capabilities.
@@
The Darcy problems, widely recognized for their applicability to multigrid methods, serve as a key focus. We anticipate substantial performance improvements compared to the current state-of-the-art. To thoroughly assess our method, we incorporated three distinct Darcy benchmarks: Darcy smooth, Darcy rough in [23], and Darcy multiscale in [24]. Darcy smooth and Darcy rough both present two-phase media within the domain, featuring different interface roughness levels. In contrast, Darcy’s multiscale presents media with multiple scales of permeability. Darcy rough and multiscale Darcy challenges our method’s ability to capture operator dependencies on fine-scale features, a task that poses difficulties for existing methods. Please refer C.1 for detailed datasets description and experiments setup.
@@
— MWT [9] only supports resolution with powers of two.
@@
Note that MgNO-high-in is not directly comparable with other models. The MgNO-high-in model differs from others, utilizing a higher input resolution of $512\times 512$ in the Darcy rough case, compared to the standard $256\times 256$ used in other models. Inspired by reduced-order modeling [26, 16, 8, 27, 30], it represents high-resolution input but targets a solution in a lower dimensional space. This is achieved by adjusting the depth/level of the $\mathcal{W}_{Mg}$ that condenses the input to the desired output resolution. Our findings suggest that high-resolution inputs enhance fine-scale predictions, implying the model effectively balances computational efficiency with high fidelity. On the other hand, by incorporating appropriate finite element basis functions, MgNO, compatible with the finite element method, demonstrates the discretization invariant capability. MgNO can train on lower-resolution datasets and evaluate on higher resolutions without the necessity of high-resolution training data, thus accomplishing zero-shot super-resolution. Please see D for the detailed experiments.
@@
Figure 3: Comparison of training dynamics between MgNO, FNO and UNO. The x-axis represents the number of epochs, and the y-axis is the error in the log scale. We present both the $L^{2}$ and $H^{1}$ training and testing accuracy (errors). For full comparisons, please refer C.1
@@
In Figure 3, we present the training dynamics of different models on benchmark Darcy rough. We train all the models in the same setting. Please see C.1 for further training settings). Despite its minimal parameter count, our model exhibits rapid convergence without susceptibility to overfitting. It is noteworthy that while larger models like MWT(9.80M), UNO(16.39M), and LSM(4.81M) can achieve training errors comparable to, or even lower than, our model, their testing errors are considerably higher. This discrepancy underscores our model’s robust generalization capabilities on unseen data.
@@
We conduct experiments on the Helmholtz equation. The Helmholtz equation poses significant challenges for numerical methods, primarily due to the resonance phenomenon. Please refer to C.3 for more details about the dataset. We present the comparison in Table 3 and Figure 7.
@@
Numerous studies have explored the use of neural operators for fluid dynamics simulation [17, 23, 36, 28]. Our focus is on the 2D Navier-Stokes equation in its vorticity form defined over the unit torus, $\mathsf{T}$ (refer to Section C.2 for an in-depth discussion). The vorticity, represented by $\omega(x,t)$, is defined for $x\in\mathsf{T}$ and $t\in[0,T]$. We aim to learn the operator $\mathcal{S}:w(\cdot,0\leq t\leq 9)\rightarrow w(\cdot,10\leq t\leq T)$, which projects the vorticity up to time 9 onto the vorticity up to a later time $T$, following the benchmark set by [23]. Our experiments consider viscosities of $\nu=1e-5$ and conclude at time $T=20$.
Furthermore, we also test models on the same Navier-Stokes task with $\nu=1e-5$ but introduce an alternative training configuration labeled as ”NS2”. It is imperative to note that in the Navier Stokes $\nu=1e-5$ configuration with training tricks, all models incorporate these specialized training techniques. These modifications consistently enhance performance when compared to the original setup. The inclusion of these training tricks contributes to more stable generalization errors across various models, justifying their use in performance comparisons. Hence, both evaluation methods offer an equitable basis for contrasting the efficacy of our approach with existing baselines. A comprehensive overview of the training configurations and specific tricks can be found in Section C.2.
Table 2: Performance on Navier Stokes
Model time params(m) Pipe NS NS2
FNO2D $\bm{2.0}$ $\bm{0.94}$ 0.67 15.56 4.57
DilResNet 5.9 1.04 0.31 23.28 10.91
U-NO 56.1 30.48 1.00 17.13 2.64
MWT 6.2 9.80 0.77 15.41 3.09
LSM 24.3 4.81 0.50 15.35 12.93
MgNO $6.4$ 2.89 $\bm{0.24}$ $10.82$ $1.63$
Table 3: Performance on Helmholtz.
Model time params(m) $L^{2}(\times 10^{-2})$ $H^{1}(\times 10^{-2})$
FNO2D 5.1 1.33 1.69 9.92
UNet 7.1 17.26 3.81 23.31
DilResNet 10.8 1.03 4.34 34.21
U-NO 21.5 16.39 1.26 8.03
LSM 28.2 4.81 2.55 10.61
MgNO 18.6 5.35 $\bm{0.80}$ $\bm{4.90}$
@@
The ablation study uses NS2 as a reference example. The baseline MgNO is configured with 5 levels, 32 channel dimensions, utilizes periodic boundary conditions, and comprises 5 architectural layers. One can find the detailed configuration in Table 7. We conduct the ablation& hyperparameters study by changing one model configuration and fixing the others. In the first and second rows, our analysis evaluated the impact of the number of levels (depth) within the MgNO architecture. Given that our structure simply merges convolution with a multigrid touch, this hyperparameter examination doubles as an ablation study for the influence of neural network depth. The outcomes highlight the performance sensitivity of MgNO to its depth. This mirrors properties observed in traditional multigrid methods where deeper algorithms are requisite for effectively mitigating low-frequency component errors. The third row emphasizes the role of boundary conditions in convolution operations, underscoring the distinct advantages of our method. Fourth row indicate adding more layers improves performance, but it also increases computational costs. The MgNO-fno configuration offers a higher-level ablation. Here, we substituted the multigrid parameterized $\mathcal{W}_{i,j}$ with the spectral convolution detailed in [23]. The ablation study reaffirms the strength of the MgNO approach, especially when integrating boundary conditions. This method not only resonates with traditional multigrid properties but also encapsulates the essence of convolution-based neural operators, delivering optimal performance in comparison to alternative convolution-based configurations.
@@ Conclusion
Moreover, there are still some questions worth future exploration. As mentioned in Section 3, the parameterization of $\mathcal{W}^{\ell}$ with multi-channel inputs remains an open challenge. One possible approach to address this might involve delving into multigrid methods for PDE systems on arbitrary dimensions, aligning with the task of solving multi-channel convolutional algebraic problems. Additionally, extending the MgNO to discrete PDEs on diverse domains with general meshes seems a logical next step. This could be achieved by incorporating the algebraic multigrid structure [42], as referenced in this study. Furthermore, MgNO achieves rapid convergence without overfitting, highlighting a promising direction for future research.
@@
    Universal approximation of $F_{i}$ on $[-M,M]^{k}$ using classical shallow neural networks $\widetilde{F}_{i}(x)$ for $x\in[-M,M]^{k}$: If $\sigma:\mathbb{R}\mapsto\mathbb{R}$ is not polynomial, given by the results in [21], there is $n_{i}\in\mathbb{N}^{+}$ and
@@ Appendix B Multigrid in convolution language
Algorithm 23 $u=\mathcal{W}_{Mg}(f;J,\nu_{\ell},n)$
1:  Input: discretized parameter function $f\in\mathbb{R}^{d\times d\times c_{f}}$, number of grids $J$, number of smoothing iterations $\nu_{\ell}$ for $\ell=1:J$, number of channels $n$ on each grid.
2:  Initialization: $f^{1}=K^{0}\ast f\in\mathbb{R}^{d_{1}\times d_{1}\times n}$, $u^{1,0}=0$, and discretized spatial size $d_{\ell}=\frac{d}{2^{\ell-1}}$ for $\ell=1:J$.
@@
Appendix C Further details on benchmarks
Benchmarks Time dependent Regular domain # dimension Boundary condition High resolution
Darcy smooth No Yes 2D Dirichlet No
Darcy rough No Yes 2D Dirichlet Yes
Darcy multiscale No Yes 2D Dirichlet Yes
Navier Stokes Yes Yes 2D Periodic No
Pipe No No 2D Mixed No
Helmholtz No Yes 2D Neumann No
@@ C.1 Darcy
The two-phase coefficients and solutions (referred to as Darcy smooth and Darcy rough in Table 1) are generated according to https://github.com/zongyi-li/fourier_neural_operator/tree/master/data_generation, and used as an operator learning benchmark in [23, 9, 2]. The coefficients $a(x)$ are generated according to $a\sim\mu:=\psi_{\#}\mathcal{N}\left(0,(-\Delta+cI)^{-2}\right)$ with zero Neumann boundary conditions on the Laplacian. The mapping $\psi:\mathbb{R}\rightarrow\mathbb{R}$ takes the value $a_{\max}$ on the positive part of the real line and $a_{\min}$ on the negative part. The push-forward is defined in a pointwise manner. The forcing term is fixed as $f(x)\equiv 1$. Solutions $u$ are obtained by using a second-order finite difference scheme on a suitable grid. The parameters $a_{\max}$ and $a_{\min}$ can control the contrast of the coefficient. The parameter $c$ controls the roughness (oscillation) of the coefficient; a larger $c$ results in a coefficient with rougher two-phase interfaces, as shown in Figure 2.
@@
We examine equation 25 that features a multiscale trigonometric coefficient adapted from [29]. This corresponds to the Darcy multiscale benchmark highlighted in Table 1. The domain, denoted as $D$, spans the area $[-1,1]^{2}$. The coefficient $a(x)$ is formulated as

### v2

@@ 5 Experiments
To substantiate the superiority of our method, we conducted comprehensive method comparisons across multiple benchmark scenarios. These benchmarks were categorized based on their association with specific partial differential equations (PDEs, namely Darcy, Navier-Stokes, and Helmholtz benchmarks. Furthermore, we incorporated tasks encompassing both regular and irregular domains, each subject to a variety of boundary conditions. This comprehensive comparison allows us to thoroughly assess the robustness and adaptability of our method, ensuring its effectiveness in real-world scenarios. For clearness, we summarize the benchmarks in Table 5. Please also refer to Section D for more detailed descriptions.
@@
We perform a comprehensive comparison with existing methods: (1) We include the original FNO (Li et al., 2020); (2) UNet (Ronneberger et al., 2015) and U-NO (Rahman et al., 2022); (3) multiwavelet neural operator (MWT) (Gupta et al., 2021); (4) Galerkin transformer (GT) (Cao, 2021); (5) latent spectral models (LSM) (Wu et al., 2023); (6) learned simulator for turbulence (DilResNet) (Stachenfeld et al., 2021). (7) Fine-tuned FNO (sFNO-v2) (Benitez et al., 2023). The baseline models were implemented using their official implementations. For consistency, we executed the baselines with default hyperparameters unless the experiments’ details specified otherwise. In cases where specific experiment details were lacking in the literature, we employed a reasonable network scale to ensure a fair comparison. In our training setup, we implemented a comprehensive approach. We trained the baselines using multiple configurations, including their default training settings encompassing different loss functions, training algorithms, and schedulers. We then reported the best results from these trials to mitigate any training-related variations. Notably, our observations revealed that all models benefited from the scheduler utilizing cosine annealing learning rates. Figure 3 will visually demonstrate that all models underwent sufficient training, ensuring that their performance accurately reflects their capabilities.
@@
The Darcy problems, widely recognized for their applicability to multigrid methods, serve as a key focus. We anticipate substantial performance improvements compared to the current state-of-the-art. To thoroughly assess our method, we incorporated three distinct Darcy benchmarks: Darcy smooth, Darcy rough in Li et al. (2020), and Darcy multiscale in Liu et al. (2022). Darcy smooth and Darcy rough both present two-phase media within the domain, featuring different interface roughness levels. In contrast, Darcy’s multiscale presents media with multiple scales of permeability. Darcy rough and multiscale Darcy challenges our method’s ability to capture operator dependencies on fine-scale features, a task that poses difficulties for existing methods. Please refer D.1 for detailed datasets description and experiments setup.
@@
— MWT (Gupta et al., 2021) only supports resolution with powers of two.
@@
Note that MgNO-high-in is not directly comparable with other models. The MgNO-high-in model differs from others, utilizing a higher input resolution of $512\times 512$ in the Darcy rough case, compared to the standard $256\times 256$ used in other models. Inspired by reduced-order modeling (Lucia et al., 2004; Hou et al., 1999; Engquist & Souganidis, 2008; Målqvist & Peterseim, 2014; Owhadi & Zhang, 2007), it represents high-resolution input but targets a solution in a lower dimensional space. This is achieved by adjusting the depth/level of the $\mathcal{W}_{Mg}$ that condenses the input to the desired output resolution. Our findings suggest that high-resolution inputs enhance fine-scale predictions, implying the model effectively balances computational efficiency with high fidelity. On the other hand, by incorporating appropriate finite element basis functions, MgNO, compatible with the finite element method, demonstrates the discretization invariant capability. MgNO can train on lower-resolution datasets and evaluate on higher resolutions without the necessity of high-resolution training data, thus accomplishing zero-shot super-resolution. Please see E for the detailed experiments.
@@
Figure 3: Comparison of training dynamics between MgNO, FNO and UNO. The x-axis represents the number of epochs, and the y-axis is the error in the log scale. We present both the $L^{2}$ and $H^{1}$ training and testing accuracy (errors). For full comparisons, please refer D.1
@@
In Figure 3, we present the training dynamics of different models on benchmark Darcy rough. We train all the models in the same setting. Please see D.1 for further training settings). Despite its minimal parameter count, our model exhibits rapid convergence without susceptibility to overfitting. It is noteworthy that while larger models like MWT(9.80M), UNO(16.39M), and LSM(4.81M) can achieve training errors comparable to, or even lower than, our model, their testing errors are considerably higher. This discrepancy underscores our model’s robust generalization capabilities on unseen data.
@@
We conduct experiments on the Helmholtz equation. The Helmholtz equation poses significant challenges for numerical methods, primarily due to the resonance phenomenon. Please refer to D.3 for more details about the dataset. We present the comparison in Table 3 and Figure 8.
@@
Numerous studies have explored the use of neural operators for fluid dynamics simulation Kochkov et al. (2021); Li et al. (2020); Stachenfeld et al. (2021); Mi et al. (2023). Our focus is on the 2D Navier-Stokes equation in its vorticity form defined over the unit torus, $\mathsf{T}$ (refer to Section D.2 for an in-depth discussion). The vorticity, represented by $\omega(x,t)$, is defined for $x\in\mathsf{T}$ and $t\in[0,T]$. We aim to learn the operator $\mathcal{S}:w(\cdot,0\leq t\leq 9)\rightarrow w(\cdot,10\leq t\leq T)$, which projects the vorticity up to time 9 onto the vorticity up to a later time $T$, following the benchmark set by Li et al. (2020). Our experiments consider viscosities of $\nu=1e-5$ and conclude at time $T=20$.
Furthermore, we also test models on the same Navier-Stokes task with $\nu=1e-5$ but introduce an alternative training configuration labeled as “NS2”. It is imperative to note that in the Navier Stokes $\nu=1e-5$ configuration with training tricks, all models incorporate these specialized training techniques. These modifications consistently enhance performance when compared to the original setup. The inclusion of these training tricks contributes to more stable generalization errors across various models, justifying their use in performance comparisons. Hence, both evaluation methods offer an equitable basis for contrasting the efficacy of our approach with existing baselines. A comprehensive overview of the training configurations and specific tricks can be found in Section D.2. Table 2: Performance on Navier Stokes Model time params(m) Pipe NS NS2 FNO2D $\bm{2.0}$ $\bm{0.94}$ 0.67 15.56 4.57 UNet 7.6 17.27 0.65 19.82 10.08 DilResNet 5.9 1.04 0.31 23.28 10.91 U-NO 56.1 30.48 1.00 17.13 2.64 MWT 6.2 9.80 0.77 15.41 3.09 LSM 24.3 4.81 0.50 15.35 12.93 MgNO $6.4$ 2.89 $\bm{0.24}$ $10.82$ $1.63$ Table 3: Performance on Helmholtz Model time params(m) $L^{2}(\times 10^{-2})$ $H^{1}(\times 10^{-2})$ FNO2D 5.1 1.33 1.69 9.92 UNet 7.1 17.26 3.81 23.31 DilResNet 10.8 1.03 4.34 34.21 U-NO 21.5 16.39 1.26 8.03 sFNO-v2 30.1 12.0 1.72 10.40 LSM 28.2 4.81 2.55 10.61 MgNO 15.1 7.58 $\bm{0.71}$ $\bm{4.02}$
@@
The ablation study uses NS2 as a reference example. The baseline MgNO is configured with 5 levels, 32 channel dimensions, utilizes periodic boundary conditions, and comprises 5 architectural layers. One can find the detailed configuration in Table 7. We conduct the ablation& hyperparameters study by changing one model configuration and fixing the others. In the first and second rows, our analysis evaluated the impact of the number of levels (depth) within the MgNO architecture. Given that our structure simply merges convolution with a multigrid touch, this hyperparameter examination doubles as an ablation study for the influence of neural network depth. The outcomes highlight the performance sensitivity of MgNO to its depth. This mirrors properties observed in traditional multigrid methods where deeper algorithms are requisite for effectively mitigating low-frequency component errors. The third row emphasizes the role of boundary conditions in convolution operations, underscoring the distinct advantages of our method. Fourth row indicate adding more layers improves performance, but it also increases computational costs. The MgNO-fno configuration offers a higher-level ablation. Here, we substituted the multigrid parameterized $\mathcal{W}_{i,j}$ with the spectral convolution detailed in Li et al. (2020). The ablation study reaffirms the strength of the MgNO approach, especially when integrating boundary conditions. This method not only resonates with traditional multigrid properties but also encapsulates the essence of convolution-based neural operators, delivering optimal performance in comparison to alternative convolution-based configurations.
@@ Conclusion
Moreover, there are still some questions worth future exploration. Our work does not sufficiently benchmark tasks on irregular domains. For irregular domains, our current implementation of MgNO requires a deformation mapping to transform the data into a regular format as shown in the pipe task. Therefore, incorporating algebraic multigrid methods Xu & Zikatanov (2017) will enhance the MgNO’s applicability to inputs and outputs in general format such as arbitrary sampling data points and point clouds. Furthermore, MgNO achieves rapid convergence without overfitting, highlighting a promising direction for future research.
Reproducibility Statement
Code Availability: The complete code used in our experiments is included in the supplementary files. Datasets can be downloaded via urls. This encompasses all scripts, functions, and auxiliary files necessary to reproduce our results. Configuration Transparency: All configurations, including hyperparameters, model architectures, and optimization settings, are explicitly provided. This ensures that every step of our experimental setup is transparent and can be replicated by others. Consistency of Experiments: To account for variability and ensure robustness in our results, each experiment was conducted a minimum of three times. Different runs were initialized with random seeds to ensure the reliability of our findings.
@@
    Universal approximation of $F_{i}$ on $[-M,M]^{k}$ using classical shallow neural networks $\widetilde{F}_{i}(x)$ for $x\in[-M,M]^{k}$: If $\sigma:\mathbb{R}\mapsto\mathbb{R}$ is not polynomial, given by the results in Leshno et al. (1993), there is $n_{i}\in\mathbb{N}^{+}$ and
@@ Appendix B Multigrid in convolution language
Algorithm 23 $u=\mathcal{W}_{Mg}(f,u^{1,0};J,\nu_{\ell},n)$
1:  Input: discretized input function $f\in\mathbb{R}^{d\times d\times c_{f}}$, discretized initial output function $u^{1,0}\in\mathbb{R}^{d_{1}\times d_{1}\times n}$, number of grids $J$, number of smoothing iterations $\nu_{\ell}$ for $\ell=1:J$, number of channels $n$ on each grid.
2:  Initialization: $f^{1}=K^{0}\ast f\in\mathbb{R}^{d_{1}\times d_{1}\times n}$, $u^{1,0}\in\mathbb{R}^{d_{1}\times d_{1}\times n}$ , $u^{1,0}=0$ if $u^{1,0}$ is not given as input, and discretized spatial size $d_{\ell}=\frac{d}{2^{\ell-1}}$ for $\ell=1:J$.
@@
Appendix C A Numerical Example Illustrating the Approximation Property of $\mathcal{W}_{Mg}$ with Sepcific Boundary Condition in equation 9
We consider elliptic PDEs described by $-\Delta u(x)=f(x)$ over the domain $\Omega=(0,1)^{2}$, subject to Dirichlet boundary conditions where $u=0$ on $\partial\Omega$. Using a linear Finite Element Method (FEM) discretization with mesh size $h=\frac{1}{(d+1)}$, the discretized equation is represented as:
$A\ast u=f,$ (25)
where $u,f\in\mathbb{R}^{d\times d}$ and $A=\left(\begin{array}[]{ccc}0&-1&0\\ -1&4&-1\\ 0&-1&0\end{array}\right)$. Here, $\ast$ denotes the standard convolution operation for a single channel, incorporating a zero-padding scheme with padding size 1 corresponding to the Dirichlet boundary condition.
In the one-channel $\mathcal{W}_{Mg}$ configuration, as detailed in Algorithm 23, we need further configurations to incorporating the Dirichlet boundary conditions. The convolutional layers $A^{\ell}$ and $B^{\ell,i}$ utilize zero-padding of size 1. In contrast, the layers $R{\ell}^{\ell+1}$ and $P_{\ell+1}^{\ell}$, for $\ell=1:J$, are set up without any padding. The convolution weights are configured as follows:
$A^{\ell}=\left(\begin{array}[]{ccc}0&-1&0\\ -1&4&-1\\ 0&-1&0\end{array}\right),\,B^{\ell,i}=\left(\begin{array}[]{ccc}0&1/64&0\\ 1/64&12/64&1/64\\ 0&1/64&0\end{array}\right),$
$R_{\ell}^{\ell+1}=\left(\begin{array}[]{ccc}0&1/2&1/2\\ 1/2&1&1/2\\ 1/2&1/2&0\end{array}\right),\text{ and }P_{\ell+1}^{\ell}=\left(\begin{array}[]{ccc}0&1/2&1/2\\ 1/2&1&1/2\\ 1/2&1/2&0\end{array}\right).$
Implementing $\mathcal{W}_{Mg}$ with these specific weights, we apply it iteratively to the right-hand side $f$ for $l$ iterations, analogous to $l$ layers in the $\mathcal{W}_{Mg}$ network. We define $u^{(l+1)}:=\mathcal{W}_{Mg}(f,u^{(l)})$ with $u^{(0)}:=0$. The convergence rate observed is illustrated in Figure 4.
Figure 4: The error, quantified on a logarithmic scale, numerically demonstrates the approximation rate, which is approximately $1-\frac{1}{c}\approx 0.1$, as outlined in equation 9.
The configurations of these weights are based on fundamental facts of piecewise linear finite elements and multigrid algorithms. For further details on these configurations, one can refer to Briggs et al. (2000); Braess (2007). For Neumann boundary conditions and periodic boundary conditions, the implementation of reflect padding and periodic padding, respectively, each with a padding size of 1, aligns seamlessly with the discretization of the finite element method, facilitating straightforward application without essential difficulty. The convergence behaviors are the same.
Appendix D Further details on benchmarks
Benchmarks Time dependent Regular domain # Dimension BD High resolution Multi-channel input
Darcy smooth No Yes 2D Dirichlet No No
Darcy rough No Yes 2D Dirichlet Yes No
Darcy multiscale No Yes 2D Dirichlet Yes No
Navier Stokes Yes Yes 2D Periodic No Yes
Pipe No No 2D Mixed No Yes
Helmholtz No Yes 2D Neumann No No
@@ D.1 Darcy
The two-phase coefficients and solutions (referred to as Darcy smooth and Darcy rough in Table 1) are generated according to https://github.com/zongyi-li/fourier_neural_operator/tree/master/data_generation, and used as an operator learning benchmark in (Li et al., 2020; Gupta et al., 2021; Cao, 2021). The coefficients $a(x)$ are generated according to $a\sim\mu:=\psi_{\#}\mathcal{N}\left(0,(-\Delta+cI)^{-2}\right)$ with zero Neumann boundary conditions on the Laplacian. The mapping $\psi:\mathbb{R}\rightarrow\mathbb{R}$ takes the value $a_{\max}$ on the positive part of the real line and $a_{\min}$ on the negative part. The push-forward is defined in a pointwise manner. The forcing term is fixed as $f(x)\equiv 1$. Solutions $u$ are obtained by using a second-order finite difference scheme on a suitable grid. The parameters $a_{\max}$ and $a_{\min}$ can control the contrast of the coefficient. The parameter $c$ controls the roughness (oscillation) of the coefficient; a larger $c$ results in a coefficient with rougher two-phase interfaces, as shown in Figure 2.
@@
We examine equation 26 that features a multiscale trigonometric coefficient adapted from (Owhadi, 2017). This corresponds to the Darcy multiscale benchmark highlighted in Table 1. The domain, denoted as $D$, spans the area $[-1,1]^{2}$. The coefficient $a(x)$ is formulated as